A total of $20,000 is invested in two bonds that pay 8% and 9.5% simple intrest. The annual intrest is $1675. How much is interested in each bond?
Answer by checkley77(12844)

.095X+.08(20,000-X)=1,675
.095X+1,600-.08X=1,675
.015X=1,675-1,600
.015X=75
X=75/.015
X=5,000 INVESTED @ 9.5%
20,000-5,000=15,000 INVESTED @ 8%
PROOF:
.095*5,000+.08*15,000=1,675
475+1,200=1,675
1,675=1,675